N2 (g) + 3 H2 (g) -> 2 NH3 (g)

assume 0.210 mol N2 and 0.674 mol H2 arw present initially. after complete reaction, how many moles of ammonia are produced? how many moles of H2 remain? how many moles of N2 remain?
Answer #1

according to balanced reaction

1 mol N2 reacts with 3 moles H2

0.210 mol N2 reacts with 0.210 x 3 / 1 = 0.630 moles H2.

but we have 0.674 mol H2.

so H2 present in excess

limiting reagent = N2.

1 mol N2 gives 2 moles NH3

0.210 mol N2 gives 0.210 x 2 / 1 = 0.420 mol NH3

moles of NH3 formed = 0.420 .

moles of H2 left = 0.674 - 0.630 = 0.044 .

moles of N2 left = 0.0 moles